OpenStudy (anonymous):

Solve: 3x - 10 - x = 2x - 4 a. x = 5/2 b. x = -1/2 c. No solutions d. Many solutions

OpenStudy (aroub):

no solutions

OpenStudy (anonymous):

3x-10-x = 2x - 4 is solved like this: 3x - x - 10 = 2x - 4 (commutative property) 2x - 10 = 2x - 4 2x - 2x = -4 + 10 (by algebra) 0x = 6 This is not possible b/c division by zero is unpermitted. If this helps, please choose good answer. For free math videos, visit http://www.tutorsean.net